Skip to main content

IT Status Alerts

All Closed Alerts

Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244

4877 results found, displaying results 381 - 401.

Scheduled Service Change: Enabling End of Service Notifications to Hosts for WebEx Meetings

Start:
2017-03-20 10:00:00
Status:
Closed
Brief Description:
We will be enabling notifications in WebEx, which will appear to hosts whenever they start a meeting. Hosts will see a notice in the web browser that the WebEx service is ending in June, and that they need to transition to Zoom.
more details

Scheduled Service Change: Update SecurID servers (Managed Server Service)

Start:
2017-03-19 05:00:00
Status:
Closed
Brief Description:
Update the SecurID servers (Managed Server Service)
more details

Scheduled Service Change: PeopleSoft: Bi-Weekly Migration 03/16/2017

Start:
2017-03-16 05:00:00
Status:
Closed
Brief Description:
PeopleSoft: Bi-weekly migration of online and batch objects.
more details

Performance Issue: Brief E-list Issues March 13, 2017

Start:
2017-03-13 13:51:00
Status:
Closed
Brief Description:
Yesterday afternoon (March 13, 2017), the main e-list server was down for two short periods, 1:51 to 2pm and 2:20 to 2:21pm. The server was returned to normal service and all messages were processed and sent out.
Current Status:
There is no problem at this time. All messages were processed and sent.
more details

Scheduled Service Change: OS-update on "nodes" hosting "shares" from SFS

Start:
2017-03-12 07:00:00
Status:
Closed
Brief Description:
Upgrade OnTap operating system on the 2 nodes in Rhodes Hall used to host Shared File Services.
more details

Scheduled Service Change: Planned maintenance for Adobe Sign

Start:
2017-03-11 08:00:00
Status:
Closed
Brief Description:
Adobe has notified us of planned downtime for Adobe Sign. The service will be unavailable in the window indicated.
more details

Scheduled Service Change: MyIT & SmartIT Upgrade

Start:
2017-03-10 08:30:00
Status:
Closed
Brief Description:
The vendor is upgrading Smart IT and MyIT, bringing new features and improvements to presets, date and time display, and templates.
more details

Performance Issue: Error Message in SSL Certificate Request Form

Start:
2017-03-06 11:34:00
Status:
Closed
Brief Description:
Website administrators requesting an SSL certificate through ssl.cit.cornell.edu may receive a message on the webpage that the request failed when it in fact succeeded.
Current Status:
The issue has been fixed. Although a previous status indicated that the number of email confirmations would be reduced, a way was found to keep the notification process essentially the same for people making requests.
more details

Performance Issue: E-list Email to Some Subdomains

Start:
2017-03-10 11:46:00
Status:
Closed
Brief Description:
Email from the E-list service (Lyris) is being delayed to some subdomains hosted by CIT, such as addresses that follow a format like account@subdomain.cornell.edu, or other domains which are hosted by CIT but don't have cornell.edu in the domain name.
Current Status:
The issue was resolved on all e-list servers by 3pm.
more details

Performance Issue: Apps Missing from Outlook on the Web Launcher

Start:
2017-03-01 13:01:00
Status:
Closed
Brief Description:
Most apps appear to be missing from the Outlook on the Web App Launcher.
Current Status:
Microsoft has announced that system patching is complete and they have verified that all previously affected systems now have the new certificate applied and service has been restored.
more details

Scheduled Service Change: Email domain CORNELL.EDU will be migrated to use the New Inbound Routers

Start:
2017-03-06 09:00:00
Status:
Closed
Brief Description:
The email domain CORNELL.EDU will be migrated to use the New Inbound Routers. This is a staged rollout that was originally scheduled to start on Monday 2/20/2017 but has been RESCHEDULED to 3/6/2017.
more details

Unplanned Outage: Phillips Network Connectivity

Start:
2017-03-06 12:48:00
Status:
Closed
Brief Description:
Network switches in Phillips Hall are down. As a result users at Phillips buildings are experiencing loss of wired and Wi-Fi data network connectivity.
Current Status:
All network switches have been restored as of 1:06PM.
more details

Scheduled Service Change: EZ-Backup team will be upgrading TSM server software on production servers

Start:
2017-03-05 02:00:00
Status:
Closed
Brief Description:
EZ-Backup team will be upgrading TSM server software on production servers
more details

Scheduled Service Change: Planned Maintenance - SkillSoft Unavailable 3/4/2017

Start:
2017-03-04 12:00:00
Status:
Closed
Brief Description:
SkillSoft will be unavailable for users during this planned maintenance period by the vendor.
more details

Unplanned Outage: Wired and Wi-Fi data network outage in Agwaste, Arthropod Research building

Start:
2017-03-03 10:11:00
Status:
Closed
Brief Description:
3 Network switches (Agwaste1-3500, Herbarium1-2800, Arthropod1-3500) are down due to NYSEG power outage.
Current Status:
NYSEG restored power, Network Switches are UP (Reachable since Mar 03, 13:53:32).
more details

Scheduled Service Change: PeopleSoft: Bi-Weekly Migration 03/02/2017

Start:
2017-03-02 05:00:00
Status:
Closed
Brief Description:
Change Description: PeopleSoft: Bi-weekly migration of online and batch objects.
more details

Scheduled Service Change: DNSDB Prod Server Hardware Upgrade

Start:
2017-02-21 05:00:00
Status:
Closed
Brief Description:
DNSDB Prod Server Hardware Upgrade
more details

Scheduled Service Change: Allow opt-in of the new Box user interface

Start:
2017-02-12 08:00:00
Status:
Closed
Brief Description:
Box will begin offering users the choice to opt-in to the new user interface before it becomes mandatory in June.
more details

Performance Issue: Believed Fixed: Multiple Services Affected by Amazon AWS Issue

Start:
2017-02-28 13:58:00
Status:
Closed
Brief Description:
(Issue believed fixed) Multiple web-based services may be affected by the issues being experienced by Amazon AWS. Amazon has announced that their engineers are continuing to work to remediate the problem.
Current Status:
Amazon believes they have resolved the issue. Some affected services may take a while to switch back to normal operations from the workarounds they put in place while Amazon tried to fix the problem. Cornell staff are continuing to monitor the situation.
more details

Performance Issue: Zoom Service Disruption

Start:
2017-02-28 13:35:00
Status:
Closed
Brief Description:
There have been reports of issues with performance of the Zoom web conferencing service.
Current Status:
Amazon has announced that they believe they have fixed the problem. According to Zoom, all services with Zoom should now be operational. Meetings that used the Cloud Recording service may experience delays in processing. Zoom engineers are continuing to monitor the system.
more details
Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244